# Datasets for Fairness in Language Models: An In-Depth Survey

### v1

@@ 3.1 Overview and Comparative Analysis
Table 1 provides a structured comparison of the major counterfactual datasets across five key dimensions defined in Section 2: structure, source, linguistic coverage, bias typology, and accessibility.
@@
WinoBias Coreference Resolution Template-based Gender–Occupation English
@@
Winogender Coreference Resolution Template-based Gender–Occupation English (+ Swed.)
@@
GAP Coreference Resolution Natural text / AI-modified Gender (names, pronouns) English
@@
StereoSet Stereotype Association Crowdsourced (contexts) Gender, Race, Religion, Profession English
CrowS-Pairs Stereotype Association Crowdsourced (minimal pairs) 9 Demographic Axes English (+ Fr., Fil., Ger., etc.)
@@
RedditBias Stereotype Association Natural Reddit Text Gender, Race, Religion, Queerness English
HolisticBias Stereotype Association Participatory Templates 13 Demographic Axes (Incl. Intersectional) English (+ Partial Multi.)
@@
EEC Sentiment / Classification Template-based Gender, Race (Names, Pronouns) English
@@
Bias NLI Natural Language Inference Template-based (premise/hypothesis) Gender, Race English
@@
Table 1: Overview of Counterfactual Fairness Datasets. The column headings follow the taxonomy defined in Section 2.2. A parallel overview for prompt-based datasets appears in Section 4.1 (Table 7).
@@
Despite their diversity in design and goals, counterfactual datasets share a common methodological principle: minimal-pair or controlled variation to assess model sensitivity to identity attributes. For example, both WinoBias and CrowS-Pairs introduce demographic shifts within near-identical contexts, but differ in their target tasks—coreference vs. stereotype classification. Likewise, stereotype and classification datasets such as HolisticBias and EEC use identity-swapped inputs to test how model scores vary with respect to demographic framing.
@@
In terms of data origin, early datasets like WinoBias, Winogender, and EEC rely on templated constructions to ensure control and clarity, while newer datasets like GAP, RedditBias, and CrowS-Pairs increasingly incorporate naturalistic data, crowdsourced content, or participatory design. Each method presents a trade-off: templates allow for experimental precision but risk unnatural phrasing; crowdsourcing offers diversity but can introduce annotator bias; and natural text ensures realism but often reflects pre-existing societal bias.
Most datasets remain English-only, although efforts such as multilingual HolisticBias and translated CrowS-Pairs demonstrate the growing need for cross-linguistic fairness assessments. From an inclusivity perspective, the field has expanded from early binary gender–occupation biases (WinoBias, Winogender) toward richer taxonomies including race, religion, LGBTQ+ identity (WinoQueer), and intersectional configurations (HolisticBias).
@@
Finally, accessibility is high across the board: most datasets are publicly released under permissive licenses (e.g., MIT, Apache 2.0, CC BY). This fosters transparency and reproducibility in fairness research, although careful attention must still be paid to construction biases and evaluation assumptions when selecting datasets for downstream model assessments.
3.2 Coreference Resolution Bias Datasets
This subsection details datasets specifically designed to evaluate fairness in coreference resolution systems. These datasets typically focus on gender bias, examining whether systems rely on societal stereotypes when linking pronouns to entities.
@@ 3.2.1 WinoBias
Description. WinoBias is a benchmark dataset introduced by Zhao et al.  [62] to evaluate gender bias in coreference resolution systems. Drawing inspiration from Winograd schemas, WinoBias presents sentence templates involving occupational nouns (e.g., “nurse,” “developer”) and gendered pronouns in contexts that require disambiguation. It includes two types of instances: Type 1 examples demand semantic reasoning (e.g., “The physician hired the secretary because {he, she} was overwhelmed with clients”), while Type 2 examples can often be resolved using syntactic cues (e.g., “The secretary called the physician and told him about a new patient”). Each template is populated using one of 40 occupations derived from U.S. Department of Labor gender statistics, as shown in Table 2. The dataset is carefully balanced with equal numbers of male- and female-pronoun references, and includes both pro-stereotypical and anti-stereotypical versions to enable a systematic assessment of bias through performance asymmetries.
@@
Dataset Taxonomy. WinoBias is a counterfactual-style dataset for coreference resolution. Each sentence pair contrasts a pro-stereotypical case with an anti-stereotypical one, where gender is ideally irrelevant to the resolution task. The dataset was generated using a template-based method, with occupations selected from U.S. labor statistics. It is monolingual (English), publicly available, and targets gender bias associated with occupational stereotypes. The design adheres to a minimal-pair paradigm, which allows for controlled experimental evaluation of bias.
@@
Representativeness Bias. WinoBias exhibits a significant representativeness bias as measured by Kullback–Leibler (KL) divergence between its occupational distribution and that of the real-world U.S. labor force. Using statistics from the Bureau of Labor Statistics (BLS), we compute the divergence between the empirical distribution in WinoBias ($P_{\mathcal{D}}$) and the population reference distribution ($P_{\mathrm{BLS}}$) as:
@@
This relatively high KL divergence indicates a strong mismatch between the dataset and real-world occupation frequencies. The source of this divergence is clearly visualized in Figure 2:
@@
  • Natural Skew in BLS Data. In contrast, BLS data show a heavy-tailed distribution: teachers are the most common occupation (6.9%), followed by managers, salespersons, cashiers, and nurses (each at 5.17%), while low-frequency occupations like tailor fall below 1%.
@@
Two pie charts further highlight the contrast: WinoBias (bottom-left) presents an artificially balanced occupational landscape, while BLS (bottom-right) reflects real-world labor dynamics.
@@
Annotation Bias. Unlike many datasets that rely on human annotators to label ambiguous or subjective content, WinoBias defines its labels deterministically. Each sentence in WinoBias has an unambiguous correct antecedent for the pronoun, derived entirely from syntactic structure rather than human judgment. Consequently, there is no risk of classic annotation bias arising from annotator prejudice, label disagreement, or cultural subjectivity.
@@
[…]
Differential Metric Bias. To evaluate whether model scoring functions exhibit asymmetric behavior across gendered sentence variants, we compute differential metric bias as defined in Equation 8. Specifically, for each scoring function $f$, we measure:
$\Delta_{f}=\mathbb{E}[f(x)\mid A=\texttt{she}]-\mathbb{E}[f(x)\mid A=\texttt{he}].$ (12)
We test for statistical significance using both a paired $t$-test and a Mann–Whitney $U$ test, and quantify practical significance using Cohen’s $d$.
Table 3 summarizes the metric-wise results. Among the six metrics tested, only Toxicity shows a small but potentially meaningful effect size ($d=0.22$), though not statistically significant. Sentiment and Regard exhibit negligible differences ($\Delta_{f}\approx 0$, $d<0.02$), suggesting scoring parity. The gender-probability metrics show small-to-large effect sizes ($d=-0.17$ to $-0.73$), but again without statistical significance, likely due to high variance or sparsity.
@@
Gender Unigram $-1.37\!\times\!10^{-3}$ 0.541 0.667 $-0.73$
Gender Max $-1.93\!\times\!10^{-3}$ 0.884 0.667 $-0.17$
Gender Wavg $-4.1\!\times\!10^{-4}$ 0.762 1.000 $-0.35$
Table 3: Differential metric bias on WinoBias (pro- vs. anti-stereotypical variants). $\Delta_{f}$: groupwise score gap; $p_{\text{t}}$: paired $t$-test; $p_{U}$: Mann–Whitney $U$-test; $d$: Cohen’s $d$.
@@ 3.2.2 Winogender
Description. The Winogender Schemas dataset, introduced by Rudinger et al.  [79], is a diagnostic resource for evaluating gender bias in English coreference resolution systems. Like WinoBias, it focuses on occupational stereotypes but adopts a minimal-pair methodology using sentence triplets that differ only by pronoun gender—specifically he, she, or singular they. This design enables the systematic assessment of coreference systems’ differential behavior across gendered references, highlighting patterns of stereotype alignment or divergence.
Dataset Taxonomy. Winogender belongs to the class of coreference resolution bias datasets and employs a counterfactual minimal-pair structure modeled after Winograd schemas. The dataset was constructed using 120 validated templates, each based on occupation-participant pairs (e.g., accountant–taxpayer). Sentence variants were manually validated through crowdsourcing (Mechanical Turk) to ensure reliable human resolution judgments. The dataset is monolingual (English), though adaptations exist in other languages (e.g., SweWinogender for Swedish). Its primary bias target is binary gender, but it also incorporates the singular they, allowing limited exploration of gender neutrality. However, the authors acknowledge the limitations of this representation. The dataset is open-source and released under the MIT License on GitHub.
@@
Bias Analysis. Following the framework outlined in Section 2.3, we assess Winogender along multiple dimensions.
@@
Annotation Bias. Coreference labels in Winogender are not crowd-sourced; they are generated deterministically from the sentence templates themselves. For every minimal pair, the correct antecedent is fixed by the syntactic schema, independent of pronoun gender. Consequently, the empirical label means are equal across all gender conditions:
@@
This equality yields the annotation bias estimator of Equation 2:
@@
Because labels are rule-derived rather than human-judged, the usual sources of annotation bias—rater prejudice, inconsistent guidelines, domain knowledge gaps—are effectively eliminated. The only residual concern is the linguistic ambiguity of singular they: although the antecedent is unambiguous in the schema, some coreference engines trained on noisy data might mis-parse these cases. From a dataset standpoint, however, the annotation process itself introduces no systematic gender skew, making Winogender an exceptionally clean test bed for isolating model-induced gender bias.
@@
Although real-world statistics (e.g., BLS data) mark nurse as female-dominated and engineer as male-dominated, Winogender’s design neutralises these priors by pairing every profession equally with all pronouns. Consequently, models cannot infer gender from co-occurrence frequency alone; any biased behaviour must stem from external knowledge encoded in the model, not from correlations inside the test set. In contrast, datasets such as CrowS-Pairs or StereoSet manifest positive PMI for stereotypical pairs, making Winogender uniquely “leak-free” at the distributional level while still probing models’ stereotype-driven heuristics through minimal pairs.
[…]
  • Sentiment. The mean gap is exactly zero for the male–female comparison and $0.0003$ for the other two contrasts. Corresponding $p$-values are $p_{t}=1.00$ and $p_{U}\approx 0.97$, with $|d|<0.001$. Hence the sentiment scorer shows no detectable gender differential.
  • Regard. The largest gap is $\Delta_{\text{reg}}=0.0021$ (male–neutral), again statistically non-significant ($p_{t}=0.81$, $p_{U}=0.96$) and negligible in magnitude ($|d|<0.023$). Thus, the regard classifier is effectively gender-invariant on this dataset.
  • Toxicity. The female–neutral comparison yields $\Delta_{\text{tox}}=0.0031$; however, $p_{t}=0.45$ and $p_{U}=0.12$ indicate no significance, and the effect size remains small ($d=0.07$). Perspective API therefore displays, at most, a weak and practically marginal differential.
  • Gender-Polarity Metrics. Metrics that explicitly quantify gendered content (Gender Unigram, Gender Max, Gender WAvg) unsurprisingly produce large, highly significant gaps ($|\Delta_{f}|\geq 0.23$, $p<10^{-40}$, $|d|\geq 1.46$). These “biases” simply confirm that the pronoun swap is detected by the polarity detector and do not reflect harmful sentiment or toxicity skew.
Across mainstream fairness metrics—Sentiment, Regard, and Toxicity—Winogender exhibits no statistically significant differential metric bias. The large gaps for gender-polarity scores serve only to validate that each minimal-pair swap is recognised, reinforcing the dataset’s suitability for controlled gender-bias probes.

### v2

@@ 3.1 Overview and Comparative Analysis
Table 1 provides a structured comparison of the major constrained-form evaluation datasets across five key dimensions defined in Section 2: structure, source, linguistic coverage, bias typology, and accessibility.
@@
WinoBias (Sec. 3.2.1) Coreference and Pronoun Resolution Template-based Gender, Occupation English
@@
Winogender (Sec. 3.2.2) Coreference and Pronoun Resolution Template-based Gender, Occupation English (+ Swedish)
@@
GAP (Sec. 3.2.3) Coreference and Pronoun Resolution Natural text / AI-modified Gender (names, pronouns) English
@@
StereoSet (Sec. 3.3.1) Sentence-Likelihood Counterfactuals Crowdsourced (contexts) Gender, Race, Religion, Profession English
CrowS-Pairs (Sec. 3.3.2) Sentence-Likelihood Counterfactuals Crowdsourced (minimal pairs) 9 Demographic Axes (e.g., gender, race, religion) English (+ French, Filipino, German, etc.)
@@
RedditBias (Sec. 3.3.3) Sentence-Likelihood Counterfactuals Natural Reddit Text Gender, Race, Religion, Queerness, Political Bias English
HolisticBias (Sec. 3.3.4) Sentence-Likelihood Counterfactuals Participatory Templates 13 Demographic Axes (Including Intersectionality) English (+ Partial Multilingual)
@@
EEC (Sec. 3.4.1) Classification-based Bias Tests Template-based Gender, Race (Names, Pronouns) English
@@
Bias NLI (Sec. 3.4.2) Classification-based Bias Tests Template-based (premise/hypothesis) Gender, Race, Ethnicity English
BBQ (Sec. 3.5.1) Question Answering (Multiple Choice) Template-based (questions/answers) Gender, Race, Occupation English
UnQover (Sec. 3.5.2) Question Answering (Underspecified) Crowdsourced (questions) Gender, Religion, Race, Ethnicity English
Grep-BiasIR (Sec. 3.6.1) Information Retrieval Bias Crowdsourced (search queries) Gender, Occupation, Race English
@@
Table 1: Overview of Constrained-Form Fairness Datasets. The column headings follow the taxonomy defined in Section 2.1. A parallel overview for open-ended datasets appears in Section 4.1 (Table 7).
@@
While these datasets differ in their specific design goals, they all follow a common methodological principle: controlled variation of identity attributes to assess model sensitivity. For example, datasets like WinoBias and CrowS-Pairs introduce demographic shifts within near-identical contexts, but they differ in their evaluation focus, with coreference resolution versus stereotype classification. Similarly, datasets like HolisticBias and EEC use identity-swapped inputs to test how model outputs vary with respect to demographic framing, such as gender or race.
@@
In terms of data source, early datasets like WinoBias, Winogender, and EEC rely on templated constructions to ensure consistency and control, while newer datasets like GAP, RedditBias, and CrowS-Pairs incorporate naturalistic data from crowdsourcing, online platforms, or participatory design. Each of these methods presents trade-offs: templated data ensures experimental precision but may lead to unnatural phrasing, while crowdsourced content introduces diversity but risks annotator bias. Natural text ensures realism but often carries the pre-existing biases of the source.
Bias Typology across these datasets is crucial for understanding the nature of the biases being tested. The majority of datasets focus on gender bias, particularly in coreference resolution and sentence completion tasks. For example, WinoBias and Winogender test how models handle gendered pronouns and whether these models exhibit a preference for certain occupations based on gender. Other datasets, such as StereoSet and CrowS-Pairs, expand beyond gender to include race, religion, and profession, testing how well models handle stereotypical associations across multiple demographic axes.
Moreover, some datasets explore more complex and intersectional biases, such as HolisticBias, which incorporates 13 demographic axes, including both gender and race, and examines the interaction between these attributes in generating biased outputs. RedditBias further adds to the complexity by including queerness and political bias, thus assessing how models respond to variations not just in demographic identity but also in more socio-political dimensions.
Most of these datasets are available in English, though efforts like HolisticBias and CrowS-Pairs demonstrate the need for cross-linguistic fairness assessments. HolisticBias even includes partial multilingual support, recognizing the importance of examining fairness in non-English contexts. The field has evolved beyond simple gender bias, as early datasets primarily examined gender-occupation stereotypes, and has moved towards more complex and intersectional taxonomies that include race, religion, LGBTQ+ identity, and even more nuanced combinations of these categories.
@@
Finally, accessibility is high across these datasets: most are publicly available under permissive licenses (e.g., MIT, Apache 2.0, or CC BY), which fosters transparency and reproducibility in fairness research. However, researchers should remain cautious of potential construction biases and the specific assumptions made in the design of these datasets when selecting them for model evaluations.
3.2 Coreference and Pronoun Resolution Datasets
This subsection details datasets specifically designed to evaluate fairness in coreference resolution systems. These datasets typically focus on gender bias, examining whether systems rely on societal stereotypes when linking pronouns to entities. Coreference resolution models are tasked with determining which noun phrases refer to the same entity within a text. The datasets in this category primarily introduce minimal-pair variations to test whether a model’s performance on pronoun resolution is biased by the gender or occupation of the entities involved. Datasets such as WinoBias, Winogender, and GAP focus on pronoun resolution by switching gendered pronouns and testing whether models show a preference for stereotypical associations, such as associating males with professions like “doctor” and females with roles like “nurse”.
@@ 3.2.1 WinoBias
Description. WinoBias is a benchmark dataset introduced by Zhao et al.  [24] to evaluate gender bias in coreference resolution systems. Drawing inspiration from Winograd schemas, WinoBias presents sentence templates involving occupational nouns (e.g., “nurse,” “developer”) and gendered pronouns in contexts that require disambiguation. It includes two types of instances: Type 1 examples demand semantic reasoning (e.g., “The physician hired the secretary because {he, she} was overwhelmed with clients”), while Type 2 examples can often be resolved using syntactic cues (e.g., “The secretary called the physician and told him about a new patient”). Each template is populated using one of 40 occupations derived from U.S. Department of Labor gender statistics, as shown in Table 2. The dataset is carefully balanced with equal numbers of male-pronoun and female-pronoun references, and includes both pro-stereotypical and anti-stereotypical versions to enable a systematic assessment of bias through performance asymmetries.
@@
Dataset Taxonomy. WinoBias is classified as a constrained-form evaluation dataset within the coreference and pronoun resolution category. It employs a minimal-pair counterfactual design where each instance presents two sentences that differ only in gender pronouns, requiring models to resolve which entity the pronoun refers to. This controlled structure isolates gender bias from other confounding factors, allowing precise measurement of accuracy differences between male and female pronoun resolution across occupational contexts.
In terms of source, WinoBias employs a template-based construction method combined with external data sources, specifically U.S. labor statistics for occupation selection. This approach provides precise control over sentence variability and demographic terms, allowing for targeted fairness evaluations while maintaining experimental rigor.
The dataset exhibits monolingual linguistic coverage, focusing exclusively on English. This English-centric approach reflects the dominance of English in NLP research but limits the generalizability of fairness conclusions across diverse linguistic and cultural contexts.
Regarding bias typology, WinoBias primarily targets gender bias associated with occupational stereotypes, representing a demographic characteristic bias. The dataset specifically addresses gender-occupation associations, which constitute a well-documented form of social bias in language models.
Finally, WinoBias demonstrates high accessibility as a publicly available dataset. The resource is openly accessible to researchers without major barriers, promoting transparency and replicability in fairness research while enabling broad community participation in bias evaluation and mitigation efforts.
@@
Representativeness Bias. WinoBias exhibits a significant representativeness bias as measured by Kullback–Leibler (KL) divergence between its occupational distribution and that of the real-world U.S. labor force. Using statistics from the Bureau of Labor Statistics (BLS) 2023, we compute the divergence between the empirical distribution in WinoBias ($P_{\mathcal{D}}$) and the population reference distribution ($P_{\mathrm{BLS}}$) as:
@@
This relatively high KL divergence indicates a strong mismatch between the dataset and real-world occupation frequencies. The source of this divergence is clearly visualized in Figure 3, and the detailed occupation frequencies from BLS 2023 are provided in Appendix A.1.1.
@@
  • Natural Skew in BLS 2023 Data. In contrast, BLS 2023 data show a heavy-tailed distribution: teachers are the most common occupation (6.9%), followed by managers, salespersons, cashiers, and nurses (each at 5.17%), while low-frequency occupations like tailor fall below 1%.
@@
Two pie charts further highlight the contrast: WinoBias (bottom-left) presents an artificially balanced occupational landscape, while BLS 2023 (bottom-right) reflects real-world labor dynamics.
@@
Annotation Bias. WinoBias is immune to classic annotation bias in the usual sense of noisy or prejudiced ground-truth labels: the correct antecedent of each pronoun is fixed deterministically by the sentence’s syntactic structure, so human judgment never enters the labeling loop. Potential bias therefore arises only from the auxiliary scoring metrics that researchers may apply when analysing model outputs. Following our unified definition of annotation bias, which subsumes both gold-label and metric-induced disparities, we measure the group-conditioned score gap for each metric $f$:
$\Delta_{f}\;=\;\mathbb{E}[f(x)\mid A=\texttt{she}]\;-\;\mathbb{E}[f(x)\mid A=\texttt{he}],$ (15)
and assess significance with a paired $t$-test and a Mann–Whitney $U$ test, reporting practical magnitude via Cohen’s $d$.
Table 3 summarizes the results of annotation bias on WinoBias. For four of the six metrics (Sentiment, Regard, and two of the gender-probability variants) the score gap is essentially zero ($|\Delta_{f}|<10^{-3}$) and effect sizes are negligible ($|d|<0.02$). Toxicity shows a modest effect ($d=0.22$), but the difference is not statistically significant at the $p<0.05$ level. Taken together, these outcomes indicate that WinoBias’s carefully balanced, minimal-pair design successfully suppresses metric-level annotation bias for most fairness scores; the small residual gap in Toxicity likely reflects bias in the toxicity instrument itself rather than in the dataset.
@@
yields a positive value for stereotypical pairs (e.g. $\text{he}\!\rightarrow\!\text{engineer}$) and a negative value for anti-stereotypical ones (e.g. $\text{she}\!\rightarrow\!\text{engineer}$). Because the dataset contains an equal number of pro and anti items, $\mathbb{E}\!\bigl[\mathrm{PMI}_{\text{role}}\bigr]=0$ over the whole corpus, but the variance of Equation 10 is large, signalling that half the sentences embed gender-congruent occupational stereotypes by construction. Thus WinoBias deliberately “hides” stereotype leakage in the coreference structure rather than in raw token frequencies, making it a targeted probe of model reliance on gendered heuristics despite seemingly balanced surface statistics.
@@
Gender Unigram $1.37\!\times\!10^{-3}$ 0.541 0.667 0.73-0.73
Gender Max $1.93\!\times\!10^{-3}$ 0.884 0.667 0.17-0.17
Gender Wavg $4.1\!\times\!10^{-4}$ 0.762 1.000 0.35-0.35
Table 3: Annotation bias (automated model bias) on WinoBias (pro- vs. anti-stereotypical variants). $\Delta_{f}$: groupwise score gap; $p_{\text{t}}$: paired tt‐test; $p_{U}$: Mann–Whitney $U$-test; $d$: Cohen’s $d$.
@@ 3.2.2 Winogender
Description. The Winogender Schemas dataset, introduced by Rudinger et al.  [51], is a diagnostic resource for evaluating gender bias in English coreference resolution systems. Like WinoBias, it focuses on occupational stereotypes but adopts a minimal-pair methodology using sentence triplets that differ only by pronoun gender, specifically he, she, or singular they. This design enables the systematic assessment of coreference systems’ differential behavior across gendered references, highlighting patterns of stereotype alignment or divergence.
Dataset Taxonomy. WinoGender is classified as a constrained-form evaluation dataset within the coreference and pronoun resolution category. Unlike WinoBias, it uses sentence triplets that differ only by pronoun gender (he, she, or singular they), enabling evaluation of how coreference systems handle binary and neutral gender references. The dataset’s Winograd schema-inspired design tests whether models rely on stereotypical associations when resolving pronouns in occupational contexts, with each template validated through crowdsourcing to ensure reliable human judgments.
In terms of source, WinoGender combines template-based construction with crowdsourced validation. The dataset was built using 120 validated templates based on occupation-participant pairs, with sentence variants manually validated through Amazon Mechanical Turk to ensure reliable human resolution judgments. This hybrid approach balances experimental control with human validation.
The dataset exhibits monolingual linguistic coverage, focusing exclusively on English, though adaptations exist in other languages such as Swedish (SweWinogender). This English-centric design reflects the dominance of English in NLP research but limits cross-cultural applicability of fairness conclusions.
Regarding bias typology, WinoGender primarily targets gender bias, representing a demographic characteristic bias. While the dataset focuses on binary gender categories, it incorporates the singular they, allowing limited exploration of gender neutrality. However, the authors acknowledge the limitations of this representation in capturing the full spectrum of gender identities.
Finally, WinoGender demonstrates high accessibility as an open-source dataset released under the MIT License on GitHub. This public availability promotes transparency, collaboration, and replicability in fairness research, enabling broad community participation in bias evaluation and mitigation efforts.
@@
Bias Analysis.
@@
Annotation Bias. The gold coreference labels in Winogender are generated deterministically from sentence templates. Each minimal pair (e.g., male–female, male–neutral, female–neutral) is constructed such that the correct antecedent is fixed by the syntactic schema, and this antecedent remains invariant across pronoun changes. Consequently, the expected label values are equal across all gender conditions:
@@
which implies an annotation bias score of
@@
Because the labels are schema-derived and not crowd-annotated, the usual sources of annotation bias, such as rater subjectivity, cultural stereotypes, or inconsistent guidelines, are effectively eliminated. The only residual ambiguity concerns the use of singular they, which some coreference models trained on noisy data may misinterpret. However, from a dataset construction standpoint, Winogender is free of label-related gender skew, making it a highly controlled benchmark for fairness diagnostics.
To assess whether external scoring functions introduce unintended disparities when applied to the dataset, we evaluated six widely used metrics:
$f\in\{\text{Sentiment},\;\text{Toxicity},\;\text{Regard},\;\text{Gender Unigram},\;\text{Gender Max},\;\text{Gender WAvg}\},$ (21)
on every minimal pair (male–female, male–neutral, female–neutral). While these scoring functions are not part of the ground truth, their behavior reveals whether off-the-shelf NLP metrics introduce unintended group-dependent signals. Table 14 (Appendix A) lists $\Delta_{f}$ (Equation 7), $t$-test and Mann–Whitney $U$ $p$-values, and Cohen’s $d$.
  • Sentiment. The male–female comparison yields a mean gap of exactly $0.0000$, with $p_{t}=1.00$ and $|d|<0.001$. The other two contrasts (male–neutral and female–neutral) show $\Delta=0.0003$ with $p\approx 0.97$. Thus, no sentiment difference emerges across gendered variants.
  • Regard. The largest observed gap is $\Delta=0.0021$ (male–neutral), with $p_{t}=0.81$ and $|d|<0.023$, indicating no meaningful disparity. The regard classifier remains invariant to gender swaps in this controlled setting.
  • Toxicity. The female–neutral contrast shows a mean difference of $\Delta=0.0031$, but with $p_{t}=0.45$ and a small effect size ($d=0.07$). Hence, no significant discrepancy is introduced by the Perspective API toxicity model.
  • Gender-Polarity Metrics. The three polarity-focused metrics (Gender Unigram, Gender Max, Gender WAvg) predictably yield large and statistically significant gaps ($|\Delta|\geq 0.23$, $p<10^{-40}$, $|d|\geq 1.46$), simply reflecting the presence of gendered pronouns in the input. These results validate that each minimal-pair swap is registered by the metric but do not indicate any evaluative bias.
In summary, both the deterministic label generation and the behavior of widely used auxiliary scoring functions indicate that Winogender exhibits no annotation bias with respect to gender. The auxiliary scorers yield statistically insignificant variations across gendered instances in core fairness dimensions such as sentiment, toxicity, and regard. This reinforces Winogender’s value as a gold-standard, stereotype-free resource for probing gender bias in coreference systems and for evaluating the fairness sensitivity of downstream NLP tools.
@@
Although real-world statistics (e.g., BLS 2023 data) mark nurse as female-dominated and engineer as male-dominated, Winogender’s design neutralises these priors by pairing every profession equally with all pronouns. Consequently, models cannot infer gender from co-occurrence frequency alone; any biased behaviour must stem from external knowledge encoded in the model, not from correlations inside the test set. In contrast, datasets such as CrowS-Pairs or StereoSet manifest positive PMI for stereotypical pairs, making Winogender uniquely “leak-free” at the distributional level while still probing models’ stereotype-driven heuristics through minimal pairs.